# GRS: Generating Robotic Simulation Tasks from Real-World Images

### v1

@@ Abstract
We introduce GRS (Generating Robotic Simulation tasks), a novel system to address the challenge of real-to-sim in robotics, computer vision, and AR/VR. GRS enables the creation of digital twin simulations from single real-world RGB-D observations, complete with diverse, solvable tasks for virtual agent training. We use state-of-the-art vision-language models (VLMs) to achieve a comprehensive real-to-sim pipeline. GRS operates in three stages: 1) scene comprehension using SAM2 for object segmentation and VLMs for object description, 2) matching identified objects with simulation-ready assets, and 3) generating contextually appropriate robotic tasks. Our approach ensures simulations align with task specifications by generating test suites designed to verify adherence to the task specification. We introduce a router that iteratively refines the simulation and test code to ensure the simulation is solvable by a robot policy while remaining aligned to the task specification. Our experiments demonstrate the system’s efficacy in accurately identifying object correspondence, which allows us to generate task environments that closely match input environments, and enhance automated simulation task generation through our novel router mechanism.
I INTRODUCTION
Real-to-sim, the problem of generating a digital twin simulation from real-world observations to train virtual agents, is a crucial challenge in computer vision, robotics, and AR/VR. This problem can naturally be broken down into three steps: 1) understanding the scene and its components, 2) finding or creating assets to populate the simulation, and 3) generating tasks to train agents in the simulated scene. Real-to-sim is valuable as this process can successfully train agents capable of acting in the real world [1]. Using these simulated environments, a real robotics system can test plans or train low-level control with high diversity without the expense and safety risks associated with real-world robot activity.
The problem of generating useful simulations from real-world observations is non-trivial. The scene must be decomposed into its parts and their relations, including the objects and their parts, the spatial arrangement of the objects, and the appearance and visual properties of the objects. Several methods exist for acquiring a digital scene, such as 3D reconstruction [2], rendering and manipulating latent space [3, 4], pose estimation with shape fitting [5], and inverse rendering [6, 7], among others. Using these scene representations, it is possible to generate specific or free-form tasks in simulation to train robots [8]. Given the value of training in simulation, recent research has focused on leveraging large language models (LLMs) to generate diverse tasks [9, 10, 11] or to generate a 3D scene from a prompt [12, 13, 14].
In this work, we present our system for Generating Robotic Simulation tasks (GRS), a complete system that extracts a scene description from a single real-world observation and generates diverse, solvable tasks (see Figure 1). Our system first identifies objects in the scene using SAM2 [15]. Each segmented part is processed by a vision-language model (VLM) to provide a text description of the object. A VLM then matches these objects against a set of simulation ready assets to produce a scene description. Using the scene description, we generate tasks that a robot could accomplish in the scene, e.g., packing a snack that contains cookies and granola bars.
@@ II RELATED WORK
Simulation Creation with LLMs has been explored to automate and enhance the creation of simulations for various applications [24, 25]. GenSim [9] uses LLMs to generate robotic simulation tasks, demonstrating the potential of LLMs to create diverse and complex simulation scenarios. Zeng et al. [10] present a system to generate reward functions based on a task definition. RoboGen [12] presents a generative robotic agent that automatically learns diverse robotic skills at scale via generative simulation. This system leverages foundation and generative models to create diverse tasks, scenes, and training supervision. Similar to our method, the system generates tasks using the scene information and also proposes a training approach based on the type of task being generated. Holodeck [14] offers language-guided generation of 3D embodied AI environments. This approach shows the potential of integrating natural language processing with 3D environment creation for robotic simulations. Compare to our method, these methods do not ensure the tasks are representative of the task definition, or scenes they generate are solvable or functional.
III METHOD
@@
Using this scene data, we formulate a task for a robotic system to execute. The extracted 3D assets and scene information are the key inputs for generating an initial simulation and associated test cases. We introduce a novel iterative refinement process, termed router, which progressively enhances both the simulation and test cases until a policy successfully completes the prescribed task. Following GenSim [9], we differentiate a task, which refers to a text description of goals and/or actions to be executed by a robotic system, and a simulation, which refers to the code our system generates to simulate that task. This distinction separates conceptual instructions and their concrete implementation in our framework.
III-A Scene Comprehension
@@
We perform scene description in a two-stage approach that combines image segmentation with image description, as illustrated in the first entry in Figure 2. This process ensures a detailed understanding of the scene, facilitating accurate simulation and task generation.
Image Segmentation. In the initial phase, we use SAM2 (Segment Anything Model 2) [15] to comprehensively segment the input image. This state-of-the-art model excels in identifying and delineating individual elements within a scene. The process often results in over-segmentation, where individual components of larger objects, such as parts of a robot, or elements of the background, are identified as separate segments. While this level of detail may seem excessive, it provides a granular foundation for our subsequent analysis, enabling a more nuanced understanding of the scene’s composition.
@@
Object Correspondence. The object correspondence process aims to link candidate objects with appropriate 3D assets for simulation. Our approach involves three steps: 1) Asset Database Creation: we create a database of 3D asset descriptions by prompting a VLM to analyze multiple renders of each asset. This process generates rich, multi-perspective descriptions of each 3D object in our asset library. 2) Candidate Object Description: we use the same VLM to describe the candidate object crops derived from our segmentation process. This description is based solely on their visual information, ensuring a consistent basis of comparison with the asset database. 3) Description Comparison: we leverage the VLM to compare these descriptions, using both the previous text description and the cropped real image. This matches each candidate object to a 3D asset or identifies it as a non-object. This step ensures that only relevant objects are considered for the simulation.
@@
III-B Simulation Generation and Evaluation
The challenge of simulation generation lies in translating real-world objectives into a simulator-compatible program for robot execution. This code must precisely define the simulator’s starting configuration and desired end state, while using tests to confirm these conditions. Crucially, the generated simulation should be optimized for feasibility, allowing a robot policy to complete it successfully within an acceptable time frame.
GRS’s simulation generation process takes as input a scene image and a scene description, summarized in Algorithm 1. Inspired by GenSim [9], we divide the simulation generation into two phases: 1) developing an abstract task definition and selecting appropriate assets, and 2) writing the simulation program for the task. Our approach enhances both steps by incorporating scene images and using a VLM for input processing, departing from GenSim’s LLM-based method. Unlike GenSim, we do not use predefined assets; instead, we leverage candidates identified during object correspondence. This allows our task generation to benefit from both the visual context of the scene and the textual descriptions of available assets.
Task Definition Generation. Given the scene descriptions and the selected visual assets our system first generates a task definition, depicted in the second entry of Figure 2. We provide the scene information, image, and asset descriptions, prompting the creation of a contextually relevant robotics task. To accommodate a variety of potential tasks we allow for the task to use a subset of the observed assets. Our focus is on generating tasks that are both practical and challenging for robotic systems. These tasks typically involve manipulating objects within the scene in specific ways, such as stacking particular items or grouping objects by category. For instance, the system might generate tasks like: “stack all red cubes on top of the blue cylinder” or “organize objects by size from left to right on the table.”
@@ III-B Simulation Generation and Evaluation
Simulation Program Generation.
@@
We next generate the simulation code using a VLM that is provided the scene image, task definition, and asset descriptions, shown in the third entry in Figure 2. To ensure the generated simulation is valid for the robot task, we also generate a test program consisting of a battery of tests. The test program is generated by providing the simulation program and the task description as input to an LLM.
@@
To align the task description and the generated simulation, we introduce a novel LLM router system that dynamically iterates over the simulation program and tests. The algorithm follows a straightforward yet powerful approach: 1) Run Tests: Execute tests on the simulation program and collect any errors. 2) Router: Use the task description and error information to determine whether to update the generated test program or the simulation program. 3) Fix: Revise the appropriate components using either a VLM for simulation code or an LLM for test code, considering inputs such as scene image, errors, and task definition. 4) Repeat this cycle until execution occurs without errors. This algorithm is visually illustrated as the last entry in Figure 2. Despite its simplicity, this process demonstrates remarkable efficacy, enabling the system to reason about multiple components and their interrelationships. By refining both the simulation and its associated tests using the task definition as guidance, our router ensures alignment between the conceptual task description and its practical implementation in the simulated environment.
@@
We focus on robotic simulations suitable for policy execution or training. To this end, we prompt the test generation process to ensure task completion by an oracle agent. The prompt includes API information for calling an oracle agent in the simulator, along with a simplified execution loop for environment observation and action. Successful execution by an oracle agent is a stringent but valuable criterion, requiring error-free code that specifies achievable objectives within the simulator’s physical constraints. While an alternative approach could involve unit tests that only check scene definition validity, we opt for testing with the oracle robot policy. This choice, although potentially incurring higher generation costs, increases the likelihood of successful downstream task generation. By using an LLM to author the tests we ensure the oracle behavior is supplemented with tests that are aligned to the task details and scene.
IV EXPERIMENTS
@@
Our proposed methodology aims to enhance accuracy in asset retrieval and improve task-simulation alignment. Given the scarcity of benchmarks for our specific area of interest—real-to-sim translation—we introduce a novel table-top robotics-inspired task. We captured 10 distinct scenes, containing an average of 15 objects each. The objects come from the HOPE dataset [26], comprising common grocery items with readily available 3D models suitable for physical simulation. We also added colored cubes and containers.
For each scene, we recorded a 1080p RGB image along with a point cloud using a ZED 2 camera. Every object is associated with its mask, 2D and 3D bounding boxes, and a textual description. Figure 3 illustrates examples from our dataset. Using this dataset, we conduct a thorough evaluation of our ability to generate appropriate correspondences. Our experiments show that using Vision Language Models (VLMs) with textual descriptions yields the highest accuracy. Further, we compare GRS’s simulation generation capabilities against previously proposed methods and qualitatively review the generation results, showing more efficiency and improved performance at generating simulations for robot execution.
IV-A Object Correspondence
@@
We designed an object correspondence experiment to assess a model’s ability to retrieve the correct assets for a captured scene. For each 3D model in our dataset, we generate three views by randomly positioning the camera around the object while maintaining focus on it. We then have a VLM produce a detailed object description based on these renders, encompassing features like shape, colors, branding, or patterns. We also produce VLM-generated text descriptions for each cropped real-world image, see Section III-A for further details.
@@
This setup allows us to evaluate various scenarios: 1) matching image text descriptions to asset descriptions (text), 2) matching images to asset descriptions (image), and 3) matching both images and text descriptions with asset descriptions (ours). For each method, we tested both GPT4o [27] and Claude-3.5-Sonnet [28].
@@
Table I presents the object retrieval results. We use F1 scores as a balanced measure of precision and recall (higher scores indicate better performance), aggregating over 10 runs of each task, where each run produced asset descriptions and text descriptions (where appropriate) before performing object correspondence. GPT4o demonstrates superior performance in all tasks.
@@
Additionally, we include a baseline using CLIP embedding distance [29]. We embed the images of the assets in the database and compare the average of the embedding of the three images of an asset to an embedding of the cropped image to find the matching asset with the smallest CLIP embedding distance. If a real-world image crop has a CLIP similarity score below 0.5 with all assets in the scene, it is considered “not an object.”
@@
Our approach outperforms both baselines as simply matching assets based on CLIP embedding distance is not robust to occlusions and variations in object poses and lighting conditions. We performed Kruskal-Wallis testing for significance since the underlying data violated the normality assumption that would be required for ANOVA testing. The Kruskal-Wallis test found statistically significant ($p<0.05$) differences for task type, model, and their interaction, indicating our results are statistically significant differences.
@@
TABLE I: Performance of different models on each task as F1 mean with 95% confidence interval (in parentheses). Highest mean in bold.
@@
IV-B Robot Task Generation

### v2

@@ Abstract
We introduce GRS (Generating Robotic Simulation tasks), a system addressing real-to-sim for robotic simulations. GRS creates digital twin simulations from single RGB-D observations with solvable tasks for virtual agent training. Using vision-language models (VLMs), our pipeline operates in three stages: 1) scene comprehension with SAM2 for segmentation and object description, 2) matching objects with simulation-ready assets, and 3) generating appropriate tasks. We ensure simulation-task alignment through generated test suites and introduce a router that iteratively refines both simulation and test code. Experiments demonstrate our system’s effectiveness in object correspondence and task environment generation through our novel router mechanism.
1 INTRODUCTION
Digital twin simulations are valuable in game generation, AR/VR, robotics, and human training simulations. The real-to-sim problem of creating digital twins from real-world observations involves three steps: 1) understanding the scene, 2) finding/creating assets to populate the scene, and 3) generating tasks for virtual agents. Beyond robotics applications, digital twins are relevant to educational games and training simulations that mirror real environments while maintaining interactive, task-oriented gameplay.
Creating simulations from real-world observations requires decomposing scenes into their components, spatial relationships, and visual properties. Existing methods include 3D reconstruction [21], manipulating latent spaces [28, 4], pose estimation [10], and inverse rendering [23, 17]. Recent research has leveraged LLMs to generate tasks [36, 41, 7] or scenes [37, 16, 38].
We present GRS, a system that extracts scene descriptions from single real-world observations and generates diverse, solvable tasks (Figure 1). Our system segments objects using SAM2 [30], processes each part with a VLM to generate text descriptions, matches these against simulation-ready assets, and creates contextually appropriate tasks for the digital twin world.
@@ 2 RELATED WORK
Simulation Creation with LLMs has been explored to automate and enhance the creation of simulations for various applications [40, 18, 32]. GenSim [36] uses LLMs to generate robotic simulation tasks, demonstrating the potential of LLMs to create diverse and complex simulation scenarios. Zeng et al. [41] present a system to generate reward functions based on a task definition. RoboGen [37] presents a generative robotic agent that automatically learns diverse robotic skills at scale via generative simulation. This system leverages foundation and generative models to create diverse tasks, scenes, and training supervision. Similar to our method, the system generates tasks using the scene information and also proposes a training approach based on the type of task being generated. Holodeck [38] offers language-guided generation of 3D embodied AI environments. This approach shows the potential of integrating natural language processing with 3D environment creation for robotic simulations. Compared to our method, these methods do not ensure the tasks are representative of the task definition, or scenes they generate are solvable or functional. FactorSim [32] uses ideas from factorized Markov Decision Processes to generate games and trains agents using reinforcement learning in those games. Our router extends the ideas in this work to improve automated testing and iteration on game code.
World models have emerged as a way to create interactive simulations. Diffusion models can be trained on data from existing games to provide interactive game engines with new controls through text or image prompting. Examples include Genie [5, 27], WHAM [22], DIAMOND [2], GameNGen [35], and Oasis [9]. Another approach is to tune video-based generative models to condition on control input for actions or camera manipulation. Examples include Promptable Game Models [24], The Matrix [14], GameGen-X [6], and GameFactory [39]. Unlike these methods we produce code readily uses existing game engines as simulators, complementing existing workflows and ensuring hard physical constraints are enforced by the simulator. These are crucial features for domains where real-world physics and behavior matter, as when games are used to train humans for real-world tasks like surgery or piloting.
3 METHOD
@@
Using this scene data, we formulate a task for a robotic system to execute. The extracted 3D assets and scene information are the key inputs for generating a simulation program and associated test cases. We introduce a novel iterative refinement process, termed router, which iteratively enhances both the simulation program and test cases until a policy successfully completes the prescribed task. Following GenSim [36], tasks are the text description of the goals and/or actions to be executed by a robotic system, and simulations are the code that implements the task. This distinction separates conceptual instructions (task) and their concrete implementation (simulation) in our framework.
While our primary focus is on robotic applications, our methodology is directly applicable to video game development. In game development, the same challenges arise: converting real environments to interactive virtual spaces with meaningful objectives. Our approach to generating solvable tasks can be used for creating game levels with appropriate difficulty, generating game mechanics tied to physical objects, and ensuring player objectives are achievable—critical aspects of game design.
3.1 Scene Comprehension
@@
Scene description follows a two-stage approach of image segmentation followed by image description, as illustrated in the first entry in Figure 2. This divide-and-conquer approach ensures a detailed understanding of the scene, facilitating accurate simulation and task generation.
Image Segmentation. We use SAM2 [30] to segment the input image into crops, which often results in over-segmentation of object parts and background elements. This granular detail provides a foundation for nuanced scene understanding. For simulation use, we map each crop to a 3D bounding box by using depth data to transform segmented pixels to 3D positions in the robot’s coordinate frame, then fitting them within axis-aligned bounding boxes.
For later use in scene generation, we map each image crop to a 3D bounding box of the candidate object. For each crop the segmented pixels are mapped to their corresponding 3D positions using depth data. The 3D coordinates are then transformed into the robot’s coordinate frame through a calibrated transformation matrix to spatially align with the robot’s environment. Once in the robot frame, the candidate object’s position and extent are fitted within an axis-aligned 3D bounding box, enabling reliable geometric matching.
@@
Object Correspondence. The object correspondence process aims to link candidate objects with appropriate 3D assets for simulation. Our approach involves three steps: 1) Asset Database Creation: We create a database of 3D asset descriptions by prompting a VLM to analyze multiple renders of each asset. This process generates rich, multi-perspective descriptions of each 3D object in our asset library. We perform this pre-processing step once and retain the text description database for reuse when evaluating different target scenes. 2) Candidate Object Description: We use the same VLM to describe the candidate object crops derived from our segmentation process. This description is based solely on their visual information, ensuring a consistent basis of comparison with the asset database. This step occurs once per real-to-sim target scene, as each scene has a different set of cropped images as output from the image segmentation. 3) Description Comparison: We use the VLM to compare the candidate object text description and the cropped real image to the descriptions in the asset database. This matches each candidate object to a 3D asset in the asset database or identifies that there is no object in the cropped image (to address over-segmentation). This step is also performed once per target scene.
@@
3.2 Simulation Generation and Evaluation
The challenge of simulation generation lies in translating real-world objectives into a simulator-compatible program for robot execution. This code must precisely define the simulator’s starting configuration and desired end state. Crucially, the generated simulation should run without errors and be optimized for feasibility, allowing a robot policy to complete it successfully within an acceptable time frame.
GRS’s simulation generation process takes as input a scene image and a scene description, summarized in Algorithm 1. Inspired by GenSim [36], we divide the simulation generation into two phases: 1) genearing an abstract task definition and selecting appropriate scene assets, and 2) writing the simulation program for the task. Our approach enhances both steps by incorporating scene images and using a VLM for input processing, departing from GenSim’s LLM-based method. Unlike GenSim, we do not use predefined assets; instead, we leverage candidate assets and their placements identified during object correspondence. This allows our task generation to benefit from both the visual context of the scene and the textual descriptions of available assets.
Task Definition Generation. Given the scene descriptions and the selected visual assets our system first generates a task definition, depicted in the second entry of Figure 2. As input we provide the scene information, image, and asset descriptions, prompting for creation of a contextually relevant robotics task. To accommodate a variety of potential tasks we allow for the task to use a subset of the observed assets. Our focus is on generating tasks that are both practical and challenging for robotic manipulation systems. These tasks typically involve manipulating objects within the scene in specific ways, such as stacking particular items or grouping objects by category. Listing LABEL:lst:task-def shows an example of a task definition generated by our system, where the LLM chose to make the task packing food items in a box.
@@ 3.2 Simulation Generation and Evaluation
Listing 1: Example task definition.
"task-name": "pack-food-items-in-box",
"task-description": "Pick up all the food items and place them inside the open box.",
"assets-used": [
"HOPE/GranolaBars.urdf",
"HOPE/MacaroniAndCheese.urdf",
"HOPE/OrangeJuice.urdf",
"HOPE/Raisins.urdf",
"HOPE/Mayo.urdf",
"HOPE/Mustard.urdf",
"HOPE/Cookies.urdf",
"container/container-template.urdf"
@@
Simulation Program Generation. We next generate the simulation code using a VLM that is provided the scene image, task definition, and asset descriptions, shown in the third entry in Figure 2. Following GenSim, the simulation code is python code that subclasses a generic Task class. The LLM prompt includes the object bounding box positions as floats and strings referencing the URDF assets to load in the simulator. The LLM is permitted to modify the object list and positions during iteration on the task.
@@
To ensure the generated simulation is valid for the robot task, we also generate a battery of tests intended to ensure the task can be completed by a robot policy. The test program is generated by providing the simulation program and the task description as input to an LLM. Tests are implemented as python unit tests that use the ‘unittest‘ library. Listing LABEL:lst:task-code in the appendix shows the final simulation output by our system for the task definition in Listing LABEL:lst:task-def.
@@
To align the task description and the generated simulation, we introduce a novel LLM router system that dynamically iterates over the simulation program and tests. The algorithm follows a straightforward yet powerful approach: 1) Run Tests: Execute tests on the simulation program and collect any errors. 2) Router: Use the task description and error information to determine whether to update the generated test program or the simulation program. 3) Fix: Revise the appropriate components using either a VLM for simulation code or an LLM for test code, considering inputs such as scene image, errors, and task definition. 4) Repeat this cycle until execution occurs without errors. This algorithm is visually illustrated as the last entry in Figure 2. This process is simple, yet effective, enabling the system to reason about the components of simulation generation and their relationships. By refining both the simulation and its associated tests using the task definition as guidance, our router ensures alignment between the conceptual task description and its practical implementation in the simulated environment. Listing LABEL:lst:test-code shows one of the tests generated by our system and Listing LABEL:lst:router shows the prompt for the router.
Listing 2: Example test code.
def test_oracle_picks_and_places_objects(self):
oracle_agent = self.task.oracle(self.env)
obs = self.env.reset()
info = self.env.info
pick_count = 0
place_count = 0
for _ in range(self.task.max_steps):
act = oracle_agent.act(obs, info)
if act is None:
break # Ensure we handle cases where the oracle cannot generate an action
obs, reward, done, info = self.env.step(act)
if act and ’pose0’ in act and ’pose1’ in act:
pick_count += 1
place_count += 1
if done:
break
# Ensure that the oracle agent attempted to pick and place objects.
self.assertGreater(pick_count, 0, "Oracle agent did not attempt to pick any objects.")
self.assertGreater(place_count, 0, "Oracle agent did not attempt to place any objects.")
Listing 3: Prompt for router.
You are an AI in robot simulation code and task design.
My goal is to design diverse and feasible tasks for tabletop manipulation.
Below are the results of running the current unit tests on the simulation code for a task.
Should I next attempt to fix the task code or fix the unit test code?
Provide a detailed explanation of what is going wrong according to the test results.
Then describe what would be the easiest fix to make to reduce the number of failing tests.
Consider whether it would be better to modify the tests to get more detailed errors or to modify the code to address the current errors.
End your reply with directions on whether to change the task code or tests.
Only provide one of the following values as a valid output, enclosed in a code block (‘‘‘):
- "fix_code" # fix the task code
- "fix_test" # fix the unit tests
Here is the task definition:
{task}
Here are the unit test results:
{test_results}
@@
We focus on robotic simulations suitable for policy execution or training. The LLM is prompted to write tests to ensure an oracle robot policy can succeed at the task. The includes API information for initializing a generic task in the simulator and calling an oracle agent in the simulator, along with a simplified execution loop for environment observation and action. Successful execution by an oracle agent is a stringent but valuable criterion, requiring error-free code that specifies achievable objectives within the simulator’s physical constraints. While an alternative approach could involve unit tests that only check scene definition validity, we opt for testing with the oracle robot policy. Testing with an oracle incurs greater simulation generation and validation costs, but increases the likelihood of successful downstream task generation. By using an LLM to author the tests using an oracle we ensure the task details represent a feasible task for downstream applications training agents in the simulator.
4 EXPERIMENTS
@@
Our method aims to improve asset retrieval accuracy and improves task-simulation alignment. Given the scarcity of benchmarks for real-to-sim translation, we introduce a table-top robotics-inspired task of 10 distinct scenes with an average of 15 objects each. Objects are from the HOPE dataset [34] of common grocery items with available 3D models, supplemented with colored cubes and containers.
For each scene, we recorded a 1080p RGB image and point cloud using a ZED 2 camera. Object correspondence generated a scene description defining objects with their crops, bounding boxes, and text descriptions. In experiments the a given robot is placed at a default center position in the scene, mirroring our scenario of converting real-world robot setup images to simulations (Figure 3).
We performed three experiments: (1) evaluating our object correspondence methodology, (2) evaluating of our task generation pipeline, and (3) demonstrating scalability to more complex scenes with larger asset databases. Results show that VLMs using images and text descriptions yield the highest accuracy for object correspondence, our simulation generation has greater efficiency and performance compared to previous methods, and our approach has potential to scale to more complex scenes with larger asset databases.
4.1 Object Correspondence
@@
Our object correspondence experiment assessed a model’s ability to retrieve the correct assets for a captured scene. For each 3D model in our dataset, we generated three rendered views of the object by randomly positioning the camera around the object in an empty space while maintaining focus on it. We then had a VLM produce a detailed object description based on these renders, encompassing features like shape, colors, branding, or patterns. A VLM also generated text descriptions for each segmented part of the real-world scene image, see Section 3.1 for further details on object description generation and object correspondence.
@@
This setup allowed us to evaluate three object correspondence methods: 1) matching image text descriptions to asset descriptions (text), 2) matching images to asset descriptions (image), and 3) matching both images and text descriptions with asset descriptions (ours). For each method, we tested both GPT4o [26] and Claude-3.5-Sonnet [3]. For object correspondence the VLM was prompted to pick the best object description from the set of generated descriptions, or indicate that no object was present.
@@
Table 1 presents the object retrieval results. We used F1 scores as a balanced measure of precision and recall (higher scores indicate better performance), aggregating over 10 runs of each task, where each run produced asset descriptions and text descriptions (where appropriate) before performing object correspondence. GPT4o demonstrated superior performance in all tasks.
@@
Additionally, we included a baseline using CLIP embedding distance [29]. We embedded the images of the assets in the database and compared the average of the embedding of the three images of an asset to an embedding of the cropped image to find the matching asset with the smallest CLIP embedding distance. If a real-world image crop had a CLIP similarity score below 0.5 with all assets in the scene, it was considered “not an object.”
@@
Our approach outperformed both baselines as simply matching assets based on CLIP embedding distance is not robust to occlusions and variations in object poses and lighting conditions. We performed Kruskal-Wallis testing for significance since the underlying data violated the normality assumption that would be required for ANOVA testing. The Kruskal-Wallis test found statistically significant ($p<0.05$) differences for task type, model, and their interaction, indicating our results are statistically significant differences.
@@
Table 1: Performance of different models on each task as F1 mean with 95% confidence interval (in parentheses). Highest mean in bold.
@@
4.2 Robot Task Generation